OpenStudy (anonymous):

Use synthetic division and the Remainder Theorem to evaluate P( c ). P(x) = x^7 - 3x^2 -1 , c=3

OpenStudy (anonymous):

very hard to write synthetic division here but i will try

OpenStudy (anonymous):

coefficients of \(p(x)=x^7-3x^2-1\) are 1 , 0 , 0 , 0, 0 , -3, 0, -1 notice you have to hold the place for all the missing terms

OpenStudy (anonymous):

Okay I see where I went wrong! Thank you :)

OpenStudy (anonymous):

1 0 0 0 0 -3 0 -1 3 3 9 27 81 243 720 2160 ____________________________ 1 3 9 27 81 240 720 2159 maybe this is right

OpenStudy (anonymous):

in this case it would be easier to evaluate directly i.e. take \[p(3)=3^7-3\times 3^2-1\]

OpenStudy (anonymous):

Yes that was correct, Thank You!

OpenStudy (anonymous):

imagine! yw